OpenStudy (unikornkiller):

A typical tip in a restaurant is 15% of the total bill. If the bill is $160, what would the typical tip be?

jhonyy9 (jhonyy9):

160 is 100% x --- 15% ------------ x= 15*160/100 = 15*1,6 = 24 x=24

jhonyy9 (jhonyy9):

hope you understand it

OpenStudy (unikornkiller):

ok thanku